Check whether the relation R ={(x, y)∈N×N | xy is the square of an integer} is an equivalence relation on N.


1
Expert's answer
2021-06-08T17:45:02-0400

Let us show that the relation "R =\{(x, y)\in\N\times \N\ |\ xy \text{ is the square of an integer}\}" is an equivalence relation on "\N".


Sinse for any "a\in\N" we have that "a\cdot a=a^2", we conclude that "(a,a)\in R", and hence the relation is reflexive.


If "(x,y)\in R," then "xy=n^2" for some "n\in\N." It follows that "yx=n^2," and thus "(y,x)\in R." Therefore, the relation "R" is symmetric.


If "(x,y)\in R" and "(y,z)\in R", then "xy=n^2" and "yz=m^2" for some "n,m\in\N." It follows that "(xy)(yz)=n^2m^2," and hence "xz=(\frac{nm}{y})^2\in\N." We conclude that "(x,z)\in R," and the relation "R" is transitive.


We conclude that the relation "R" is an equivalence relation on "\N".